OpenStudy (anonymous):

Two divers jump from the same boat. One diver descends 30 feet below the ocean's surface. He then swims 85 feet west. The other diver descends 40 feet and swims 155 feet east. What is the distance between the two divers?

OpenStudy (anonymous):

sqrt(10^2+155^2)+85

OpenStudy (anonymous):

ans in feet

OpenStudy (anonymous):



OpenStudy (anonymous):

sorry d1=sqrt(10^2+155^2) i for got sqrt

OpenStudy (anonymous):

required distance =sqrt((40-30)^2+(155+85)^2)=sqrt(100+57600)=sqrt(67600)=260